### v1

Abstract
The Axelrod approximation is widely used in astrophysical modelling codes to evaluate electron-impact excitation effective collision strengths for forbidden transitions. Approximate methods such as this are a necessity for many heavy elements with open shells where collisional data is either non existent or sparse as the use of more robust methods prove prohibitively expensive. Atomic data for such forbidden transitions are essential for producing full collisional radiative models that do not assume Local-Thermodynamic-Equilibrium (LTE). In this short work we repeat the optimization of the simple Axelrod formula for a large number of -matrix data sets, ranging from Fe and Ni to the first r-process peak elements of Sr, Y and Zr, to higher Z systems Te, W, Pt and Au. We show that the approximate treatment of forbidden transitions can be a significant source of inaccuracy in such collisional radiative models. We find a large variance of the optimized coefficients for differing systems and charge states, although some general trends can be seen based on the orbital structure of the ground-state-configurations. These trends could potentially inform better estimates for future calculations for elements where -matrix data is not available.
keywords:
[…]
Research highlights item 1
Research highlights item 2
Research highlights item 3
1 Introduction
The accurate calculation of atomic data for collision processes is computationally expensive and typically has to be carried out on a case-by-case basis for each ion and ionization stage of interest. For this reason, in models where the approximation of Local-Thermodynamic-Equilibrium (LTE) cannot be assumed, the employed codes typically make use of approximate or semi-empirical datasets. The formulae of Axelrod [1] and van Regemorter [2] are often used when calculating the rates of electron-impact-excitation. Examples of works making use of these approximations include [3, 4, 5]. While such formulae are valid for producing usable data for the simulation of astrophysical transients, it has been shown recently by Bromley et al. [7] and Mulholland et al. [8] that the data these methods produce in their current state are typically inaccurate for many transitions. Generally for forbidden lines the collision strengths are underestimated by several orders of magnitude when compared with results from the -matrix method of Burke [6]. Additionally, at low temperatures the -matrix method is more reliable due to the inclusion of resonances in the low energy interaction region. The inclusion of resonances is dependent on the specific atomic system, which increases computational complexity - however their neglection can potentially reduce the calculated rates by a significant margin.
Recently, a large amount of atomic data has become available for heavy elements of astrophysical interest [9, 8, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 7], particularly those for kilonovae (KNe), supernovae (SNe) and fusion applications. In this work we compare optically thin emission spectra calculated using the Axelrod [1] and van Regemorter [2] empirical formulae with models based on data produced by sophisticated -matrix calculations. In particular, the validity of the Axelrod formula for forbidden transitions in the context of modelling KNe is investigated. Use of the Axelrod and van Regemorter formulae has been discussed previously [7, 22, 8], and some revision has also been put forward by Pognan et al. [4]. In this publication we add to these discussions by presenting an ion-dependent re-optimization of the Axelrod formula based on the availability of the new -matrix data.
In Section 2, the theory of collisional-radiative-modelling and electron-impact-excitation is briefly introduced, as well as the approximate formulae of van Regemorter [2] and Axelrod [1]. It is also demonstrated, by way of an illustrative example, how the Axlerod formula in its unaltered form can systematically under predict emission from ions. In Section 3, we detail the re-optimization of the coefficients in the Axelrod formula, where we optimize toward newly available calibrated -matrix data. Optically thin emission spectra are constructed using a selection of datasets to show how the Axelrod formula can be amended by multiplying the coefficients by an ion-dependent scaling factor. Some basic trends in terms of the shell-behaviour of the ions considered are further analysed. In Section 4 we conclude with a summary of the main findings of the paper and an outlook.
2 Theory
Modelling codes that do not assume LTE often solve the collisional-radiative equations for the level populations of ions in the respective plasma,
where are the level populations and the rates of the included atomic processes are encapsulated by the matrix . From the steady state populations, emission can be calculated using the Einstein A-coefficients , so that
where the populations are appropriately normalised, is the electron density, and is the Photon-Emissivity-Coefficient for the transition. More details surrounding collisional radiative modelling can be found in Summers et al. [23], Burke [6].
A particularly important atomic processis electron-impact excitation. This is the process where an incident electron collides with a target atom (or ion) and energy transfer occurs, promoting or demoting the target atom between energy eigenstates. Let the lower and upper level indices be and respectively with energy levels and with . The strength of a transition between these levels is typically quantified by a collision strength . To calculate the corresponding rate to be used in Eq. (1), these are normally averaged over a Maxwellian velocity distribution, so that
where is the electron temperature in energy units, and is the final energy of the incident electron. are called the effective collision strengths. The rates to be inserted into are then given by,
where Eq. (4) and (5) correspond to the excitation and de-excitation rates respectively, and here is in Kelvin. The quantities are the statistical weights of the atomic levels and is the energy separation between the two levels. Current large scale atomic modelling, such as those presented in Shingles et al. [5] and Pognan et al. [3], typically adopt that have been generated using semi-empirical formulae, as opposed to the more accurate and more computationally expensive -matrix methods of Burke [6]. The primary reason for the use of approximations in the past has been the lack of accurate collisional atomic data.
Electric dipole transitions are often estimated using the van Regemorter [2] formula, given as
where , is an effective Gaunt factor tabulated in [2] for neutral and singly ionized species, is the relevant Einstein A-coefficient, is a statistical weight, and is the wavelength of the transition in units of cm. This formula was given in van Regemorter [2] for neutral and singly ionized species. For species more ionized than this one can make use of other approximations such as those detailed in the ADAS routine d7pyvr [24]. From [samspon92] (they use as lower weight)
Using ( is upper)
Gives:
Put and integrate both sides to arrive at eq. 6. So I conclude it should probably be the upper weight.
By contrast, and the focus of this work, forbidden transitions have in the past been calculated using the Axelrod [1] formula, so that
which depends only on the statistical weights of the atomic levels. Notably they proposed the seperate treatment of lines with wavelength larger than 10m, although we focus on this base formula in this work. This is notably independent of temperature. This is likely because the Fe iii dataset of Garstang et al. [25] used by Axelrod [1] to derive this formula contained only a single for each transition. These were then used as surrogate effective collision strengths , so that for constant . In general the values associated with forbidden lines exhibit less of a dependence on temperature than those of allowed lines. This could perhaps justify the temperature independent formula for forbidden lines given in Eq. 7.
To motivate our work, we show in Figure 1, where typical examples of effective collision strengths as a function of electron temperature are shown for a forbidden line (left panel) and a strong allowed transition (right panel) of Y ii [8]. Specifically, the values computed using the -matrix method are compared with the Axelrod approximate values for the forbidden line (5s4d 3D1 - 5s4d 3D2) and the van Regemorter values for the allowed line (5s2 1S0 - 4d5p 1P). Clearly the magnitude and temperature dependence is relatively well captured by the van Regemorter approximation for the allowed transition. We found this to be the case across many transitions for this ion (see for example [8]). By contrast the Axelrod values systematically underestimate of the forbidden line collision strength by up to several orders of magnitude. We find similar discrepancies for many other forbidden transitions in Y ii.
In this work we extend the work of Bromley et al. [7], who previously discussed how these formulae can systematically under predict optically thin emission, and add to the discussions by showing the wide variance between the Axelrod formula rates and those produced by perturbative/non-perturbative collision calculations. We further show that this approximation introduces the greatest uncertainty for diagnostic modelling based upon forbidden lines. A recent publication by Pognan et al. [4] has computed temperature and ion-stage dependent forward scaling factors for both of these empirical formula, based on collisional data presented in Bromley et al. [7] for Pt () ions. They generally find the Axelrod formula needs to be scaled up by a factor for better agreement with -matrix data, which we also find here for a large number of cases. In addition, Shingles et al. [5] presents an example where the Axelrod formula was modified by amending the scaling factor to 0.01 which further emphasises the sensitivity of the cooling to the treatment of forbidden lines. We present in the next section a comprehensive analysis where multiple ion stages of several heavy elements from Fe through to W have had the rates describing forbidden transitions systematically optimized using sophisticated -matrix collisional data and compared against those produced using the Axelrod formula.
3 Reoptimization
We reoptimise the Axelrod approximation for a particular element by fitting the formula,
with lower and upper levels and respectively, and corresponding statistical weights and . Due to recently published -matrix data for several heavy systems [9, 8, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 7], it is now possible to reoptimize this coefficient for a large number of systems, retaining Equation (8) as the formula used for forbidden transitions. In particular, we find the optimal on an ion-by-ion basis.
3.1 Methods
We define two cost functions for optimising the Axelrod coefficient as,
where the inner sum is taken over all transitions with upper and lower levels and respectively and the outer sum is taken over all temperature points. We have also included optimisation with respect to a weight matrix , where allowed transitions have identically. This is so that only forbidden transitions are included in the optimization - as allowed transitions are typically calculated in other approximate manners such as van Regemorter [2].
A forbidden transition is sometimes defined in the literature as one with an oscillator strength [5, 3], although there is a more rigorous definition based on the quantum numbers of the states involved in the transition111 Technically, allowed transitions are those with a change in angular momentum () and a parity change . A forbidden transition is then one which does not satisfy these selection rules. However, for consistency with the use of the Axelrod formula in the literature, we employ this definition based on the oscillator strength.. In optimizing against -matrix data, it was found this criterion based on solely the oscillator was too lenient, as some allowed transitions can exhibit large values with small oscillator strengths. For this reason we introduce the additional requirement that on average for a forbidden transition. We therefore identify a forbidden line based on both the condition on the oscillator strength as well as the respective -matrix values. Additionally, since the primary motivation of this work is the use of Axelrod in late-stage KNe modelling, we restrict the optimization to low temperatures. We thus set to roughly encapsulate the temperature range of concern. Both optimizations and naturally amount to fitting a straight line through the origin, although we show both to emphasise the heavy dependence of the optimization on the weighting of the included transitions.
3.2 Results
figure*[ht!]
Optically thin emission spectra of Y ii at an electron temperature of K and electron density cm-3.
The Axelrod approximation was optimised for a number of atomic species for which R-matrix data was readily available. Reoptimized coefficients are shown in Table 1. The values of are those that minimized the function in Eq. (9) and the values of are those that minimized the function in Eq. (10). Elements are grouped by the highest orbital angular momentum in their ground state configuration222Except for Y iii, which is grouped with the s-shell elements due to its effective-single-electron structure.. It can be seen that quite generally is larger than for each element, in some cases by an order of magnitude. Given that the Axelrod formula is typically used for the ad hoc representation of the general behaviour within a collisional radiative model, the larger values found in will more accurately represent the cooling. It is seen from Table 1 that the -shell elements generally require a factor 10 - 100 times larger than that of the standard Axelrod formula. Similarly, -shell elements require a factor around 10-20 times higher. By contrast, the -shell elements appear well represented in order of magnitude by the stock formula. This is perhaps expected given the Axelrod formula [1] was originally optimized on the Fe iii (3d6) data of Garstang et al. [25]. With expanding atomic data sets, it will be interesting to see the impact of -shell elements when large amounts of atomic data become available for them.
table*[!ht] Optimized Axelrod-coefficients for a selection of ion stages with R-matrix data. ToDo: {Need to cite these data sets in one way or another. Super scripts?}. Ion Config Ion Config -shell -shell cont Sr i 4p6 5s2 0.05366 0.02609 Ni ii 3d9 0.00290 0.00037 Y ii 4p6 5s2 0.04460 0.03247 Ni iii 3d8 0.00668 0.00201 Sr ii 4p6 5s 0.47730 0.21053 Ni iv 3d7 0.00472 0.00121 Y iii 4p6 4d 0.37418 0.23215 Zr i 4d2 5s2 0.00503 0.00212 Zr ii 4d2 5s 0.01674 0.00631 -shell Zr iii 4d2 0.05006 0.02053 Sr iii 4p6 0.07666 0.04206 W i 6s2 5d4 0.00283 0.00139 Te i 5s2 5p4 0.06736 0.01168 W ii 6s2 5d3 0.00372 0.00203 Te ii 5s2 5p3 0.07460 0.05840 W iii 6s2 5d2 0.00653 0.00431 Te iii 5s2 5p2 0.08868 0.06478 Pt i 5d9 6s 0.00543 0.00148 Pt ii 5d9 0.00444 0.00174 -shell Pt iii 5d8 0.00309 0.00147 Fe ii 3d6 4s 0.00233 0.00058 Au i 5d10 6s 0.01183 0.00554 Fe iii 3d6 0.00274 0.00075 Au ii 5d10 0.00471 0.00055 Au iii 5d9 0.00373 0.00176
Ion | Config | Source | Ion | Config | Source | ||||
-shell | -shell | cont | |||||||
Sr i | 4p6 5s2 | 0.05366 | 0.02609 | [9] | Ni ii | 3d9 | 0.00290 | 0.00037 | [14] |
Y ii | 4p6 5s2 | 0.04460 | 0.03247 | [8] | Ni iii | 3d8 | 0.00668 | 0.00201 | [15] |
Sr ii | 4p6 5s | 0.47730 | 0.21053 | [8] | Ni iv | 3d7 | 0.00472 | 0.00121 | [16] |
Y iii | 4p6 4d | 0.37418 | 0.23215 | [11] | Zr i | 4d2 5s2 | 0.00503 | 0.00212 | [17] |
Zr ii | 4d2 5s | 0.01674 | 0.00631 | [17] | |||||
-shell | Zr iii | 4d2 | 0.05006 | 0.02053 | [17] | ||||
Sr iii | 4p6 | 0.07666 | 0.04206 | [9] | W i | 6s2 5d4 | 0.00283 | 0.00139 | [18] |
Te i | 5s2 5p4 | 0.06736 | 0.01168 | [10] | W ii | 6s2 5d3 | 0.00372 | 0.00203 | [19] |
Te ii | 5s2 5p3 | 0.07460 | 0.05840 | [10] | W iii | 6s2 5d2 | 0.00653 | 0.00431 | [20] |
Te iii | 5s2 5p3 | 0.08868 | 0.06478 | [10] | Pt i | 5d9 6s | 0.00543 | 0.00148 | [7] |
Pt ii | 5d9 | 0.00444 | 0.00174 | [7] | |||||
-shell | Pt iii | 5d8 | 0.00309 | 0.00147 | [7] | ||||
Fe ii | 3d6 4s | 0.00233 | 0.00058 | [12] | Au i | 5d10 6s | 0.01183 | 0.00554 | [21] |
Fe iii | 3d6 | 0.00274 | 0.00075 | [13] | Au ii | 5d10 | 0.00471 | 0.00055 | [21] |
Au iii | 5d9 | 0.00373 | 0.00176 | [21] |
The emission spectrum as a function of wavelength in nm (scaled relative to the strongest line using only -matrix data) for Y ii is shown on Figure 2. The left hand panel depicts three optically thin emission spectra using different Y ii datasets. Namely, the full R-matrix data for both allowed and forbidden transitions (RM+RM), the R-matrix data for electric dipoles and the baseline Axelrod formula (=0.004) for forbidden lines (RM+AX) and finally the van Regemorter formula for allowed lines and the baseline Axelrod formula for forbidden lines (VR+AX). By contrast, the right hand panel again uses the full R-matrix data (RM+RM), the R-matrix data for electric dipoles and the modified Axelrod formula (=0.044) for forbidden lines (RM+AX) and finally the van Regemorter formula for allowed lines and the modified Axelrod coefficient for forbidden lines (VR+AX). It is clear that the emission spectra are more comparable when the modified Axelrod approximation is adopted. It is of note that the agreement increases with only a change to the approximate formula for forbidden transitions. This is an indication that the accuracy of NLTE collisional radiative models is limited by the treatment of forbidden transitions. This is perhaps not surprising, as forbidden transitions provide efficient thermal excitation/de-excitation pathways at low electron temperatures.
We additionally show on Figure 3 the scaled emission for Au i, employing a similar set of modified data to the above. Au i presents a special case of a -shell element whose electronic configurations are constructed from the 5d10 and 5d9 series. The former set behaves like a single electron system. There is therefore a limited number of forbidden transitions that are typically stronger than those of other systems. For this reason the optimized Axelrod coefficient is considerably higher than the surrounding -shell elements presented in Table 1. Figure 3 therefore demonstrates a mixed bag of agreement in the emission estimations because the Axelrod formula assumes that all the forbidden lines behave the same, which is evidently not the case for this ion.
It can be seen that certain lines are well represented with the R-matrix + Axelrod datasets (RM+AX), regardless of the Axelrod coefficient used. It was found these represent dipole allowed lines whose upper level population is dominated by the corresponding excitation rate, which are represented by the R-matrix data in that formalism. It is noteworthy that such lines are typically poorly represented by the van Regemorter + Axelrod data (VR+AX), showing the limiting factor for these lines at this point in temperature and density space are actually the effective collision strengths of allowed transitions. Two examples can be seen at nm, which are highlighted in grey on Figure 3. Notably both mixes of data show little benefit from the improved Axelrod approximation. In fact, those lines that were truly dominated by the R-matrix dipole data are in some cases over-estimated by the improved Axelrod approximation (see for example the lines between 300 and 400nm highlighted in red).
By contrast (and is the case for the majority of the lines shown) there is significant underestimation of the emission in the case of the base Axelrod approximation, where agreement is seen between the data sets employing this with either the R-matrix or van Regemorter allowed rates. Moving to the re-optimized Axelrod approximation introduces a certain improvement in the estimation of the emission, in many cases by an order of magnitude. The main victory is the improvement of the evidently strongest lines likely to contribute to a real astrophysical spectrum.
In principle we could optimise only on the strongest lines, although the optimization would depend on the plasma conditions chosen. For a low density plasma, the coronal approximation applies and the optimization would reduce to fitting Eq. (8) to the subset of effective collision strengths . For a high density plasma, LTE applies and in principle the populations and thus the emission should be independent of the rates. Given that the particular positions in density space of the coronal, collisional-radiative and thermal equilibrium regimes are vastly different for different ion stages, perhaps a general optimization over all the available rates as performed here remains the best solution as opposed to the choice of particular lines. Such an effort may be feasible when a greater breadth of -matrix data spanning the full range of , , , and shell elements relevant to KNe becomes available.
4 Conclusions and Outlook
In my opinion the conclusions are quite basic. Depsite the progress that had been made in this work, there is a relative lack of consensus regarding the value of the Axelrod coefficient to use. The fact there is such significant variance across the ion stages shows that while the values here can be adopted, it is more correct to use the available -matrix data. The values here could in principle be used to generate collision strengths for collisions between high-lying levels not included in structure calculations. Some minor trends can be seen. Elements with -shell ground states tend to have transitions 10 - 1000 times stronger than the base Axelrod approximation. -shell ground states tend to be 10-20 times stronger. -shell elements tend to have forbidden transitions of the rough order of magnitude of the original Axelrod approximation. Despite this, the relatively large variance indicates that the Axelrod approximation should be reserved for where no data is available, and should be replaced with -matrix (or even distorted wave data, [7])
To summarise, the Axelrod approximation for effective collision strengths has been reviewed and revisited. The Axelrod coefficient has been calculated for a variety of ion stages with available R-matrix data. The elements studied cover the 1st (Sr-Zr), 2nd (Te) and 3rd (W,Pt,Au) peaks of the rapid nucleosynthesis processes. The values presented here could in principle be used to generate collision strengths for collisions between high-lying levels not included in structure calculations, or aid in the calculation of similarly structured elements by using ball-park coefficients reflective of those presented in Table 1. Elements with -shell ground states tend to have transitions 10 - 1000 times stronger than the base Axelrod approximation. -shell ground states tend to be 10-20 times stronger. -shell elements tend to have forbidden transitions of the rough order of magnitude of the original Axelrod approximation. Despite this, the relatively large variance in indicates that the Axelrod approximation should be reserved for where no data is available, and should be replaced with -matrix data as it comes available. Nonetheless, it is intended that this work serves as an intermediate step in the improvement of atomic data for astrophysical models, where the values presented here can be used to provide more information for elements with no data available and the Axelrod formula can be scaled based on the electronic configurations of the species involved.
CRediT authorship contribution statement
Leo P. Mulholland: Writing – review & editing, Writing – original draft, Visualization, Validation, Methodology, Investigation, Conceptualization, Formal analysis.
Steven J. Bromley: Writing – review & editing, Writing – original draft, Visualization, Validation, Methodology, Investigation, Conceptualization, Formal analysis.
Connor P. Ballance: Writing – review & editing, Writing – original draft, Validation, Supervision.
Stuart A. Sim: Writing – review & editing, Writing – original draft, Validation, Methodology, Visualization, Supervision.
Catherine A. Ramsbottom: Writing – review & editing, Writing – original draft, Validation, Methodology, Visualization, Supervision.
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Data Availability
The relevant atomic data can be found in the cited references.
Acknowledgements
We thank our colleagues at Queen’s University Belfast and Auburn University for helpful discussion. LPM, CPB, SAS and CAR acknowledge funding from the European Union (ERC, HEAVYMETAL, 101071865). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. SJB acknowledges NASA APRA for funding part of this work under grant number is 23-APRA23-0086.

### v2

@@
 Atomic data for such forbidden transitions are essential for producing full collisional radiative models that do not assume Local-Thermodynamic-Equilibrium (LTE).
-In this short work we repeat the optimization of the simple Axelrod formula for a large number of -matrix data sets, ranging from Fe and Ni to the first r-process peak elements of Sr, Y and Zr, to higher Z systems Te, W, Pt and Au.
+In this short work we present the re-optimization the simple Axelrod formula for a large number of -matrix data sets, ranging from Fe and Ni to the first r-process peak elements of Sr, Y and Zr, to higher systems Te, W, Pt and Au.
 We show that the approximate treatment of forbidden transitions can be a significant source of inaccuracy in such collisional radiative models.
@@
 Generally for forbidden lines the collision strengths are underestimated by several orders of magnitude when compared with results from the -matrix method of Burke [6].
-Additionally, at low temperatures the -matrix method is more reliable due to the inclusion of resonances in the low energy interaction region.
+Additionally, at low temperatures the -matrix method is more reliable when compared with methods such as distorted-wave due to the inclusion of resonances in the low energy interaction region.
 The inclusion of resonances is dependent on the specific atomic system, which increases computational complexity - however their neglection can potentially reduce the calculated rates by a significant margin.
@@
 [4].
-In this publication we add to these discussions by presenting an ion-dependent re-optimization of the Axelrod formula based on the availability of the new -matrix data.
+In this publication we add to these discussions by presenting an ion-dependent re-optimization of the numerical constant within the Axelrod formula based on the availability of the new -matrix data.
+The work here serves to illustrate the degree in variation in characteristic collision strengths between different ions, and quantify the extent to which results from full calculations can be captured with such simple formulae.
+We also discuss how existing data can serve to guide better informed approximation for elements where no -matrix data is available.
 In Section 3, the theory of collisional-radiative-modelling and electron-impact-excitation is briefly introduced, as well as the approximate formulae of van Regemorter [2] and Axelrod [1].
-It is also demonstrated, by way of an illustrative example, how the Axlerod formula in its unaltered form can systematically under predict emission from ions.
+It is also demonstrated, by way of illustrative examples, how the Axelrod formula in its unaltered form can systematically under predict emission from ions.
 In Section 4, we detail the re-optimization of the coefficients in the Axelrod formula, where we optimize toward newly available calibrated -matrix data.
-Optically thin emission spectra are constructed using a selection of datasets to show how the Axelrod formula can be amended by multiplying the coefficients by an ion-dependent scaling factor.
+Optically thin emission spectra are constructed using a selection of datasets to show how the Axelrod formula can be amended with an ion-dependent coefficient.
 Some basic trends in terms of the shell-behaviour of the ions considered are further analysed.
 In Section 5 we conclude with a summary of the main findings of the paper and an outlook.
+2 Theory
+Collisional radiative modelling codes solve the rate equations for the level populations of ions in a plasma,
+where the matrix encapsulates the rates of the atomic processes involved. This includes the electron-impact excitation and de-excitation rates given by
+where is the electron temperature in Kelvin, is the Boltzmann constant, is the statistical weight of the lower (upper) level and is the energy difference between the levels.
+The quantities represent Maxwellian-averaged-collision strengths.
+These have been calculated a number of ways, including distorted wave [23] and the R-matrix codes [6].
+As discussed in the introduction, in lieu of data produced by these methods, semi-empirical formulae are employed, such as those presented in [2, 1].
+The focus of this work is on forbidden lines, for which modelling codes typically employ the formula of [1],
+which was optimized on data from [24].
+This coefficient of 0.004 was specifically produced for data of Fe, and will be shown to result in a systematic underestimation of atomic line emission.
+The presented formula is independent of temperature, which is discussed in [25].
 3 Theory
@@
 where the populations are appropriately normalised, is the electron density, and is the Photon-Emissivity-Coefficient for the transition.
-More details surrounding collisional radiative modelling can be found in Summers et al.
-[23], Burke [6].
+More details surrounding collisional radiative modelling can be found in Summers2006, Burke [6].
 A particularly important atomic processis electron-impact excitation.
@@
 This formula was given in van Regemorter [2] for neutral and singly ionized species.
-For species more ionized than this one can make use of other approximations such as those detailed in the ADAS routine d7pyvr [24].
-From [samspon92] (they use as lower weight)
-Using ( is upper)
-Gives:
-Put and integrate both sides to arrive at eq. 6. So I conclude it should probably be the upper weight.
+For species more ionized than this one can make use of other approximations such as those detailed in the ADAS routine d7pyvr [openadas_manual].
 By contrast, and the focus of this work, forbidden transitions have in the past been calculated using the Axelrod [1] formula, so that
@@
 This could perhaps justify the temperature independent formula for forbidden lines given in Eq.
-7.
+11.
 To motivate our work, we show in Figure 1, where typical examples of effective collision strengths as a function of electron temperature are shown for a forbidden line (left panel) and a strong allowed transition (right panel) of Y ii [8].
@@
 (14).
-Elements are grouped by the highest orbital angular momentum in their ground state configuration222Except for Y iii, which is grouped with the s-shell elements due to its effective-single-electron structure..
+Elements are grouped by the highest orbital angular momentum in their ground state configuration.
 It can be seen that quite generally is larger than for each element, in some cases by an order of magnitude.
@@
 With expanding atomic data sets, it will be interesting to see the impact of -shell elements when large amounts of atomic data become available for them.
+Additionally, we have presented the mean values of within the table, for a more universal replacement in models for ions where accurate data is not yet available.
+Use of these mean values should come with caution, as there remains large spread within the shell groups.
 table*[!ht] Optimized Axelrod-coefficients for a selection of ion stages with R-matrix data.
@@
 | Ion | Config | Source | Ion | Config | Source | ||||
--shell | -shell | cont | |||||||
-Sr i | 4p6 5s2 | 0.05366 | 0.02609 | [9] | Ni ii | 3d9 | 0.00290 | 0.00037 | [14] |
-Y ii | 4p6 5s2 | 0.04460 | 0.03247 | [8] | Ni iii | 3d8 | 0.00668 | 0.00201 | [15] |
-Sr ii | 4p6 5s | 0.47730 | 0.21053 | [8] | Ni iv | 3d7 | 0.00472 | 0.00121 | [16] |
-Y iii | 4p6 4d | 0.37418 | 0.23215 | [11] | Zr i | 4d2 5s2 | 0.00503 | 0.00212 | [17] |
-Zr ii | 4d2 5s | 0.01674 | 0.00631 | [17] | |||||
--shell | Zr iii | 4d2 | 0.05006 | 0.02053 | [17] | ||||
-Sr iii | 4p6 | 0.07666 | 0.04206 | [9] | W i | 6s2 5d4 | 0.00283 | 0.00139 | [18] |
-Te i | 5s2 5p4 | 0.06736 | 0.01168 | [10] | W ii | 6s2 5d3 | 0.00372 | 0.00203 | [19] |
-Te ii | 5s2 5p3 | 0.07460 | 0.05840 | [10] | W iii | 6s2 5d2 | 0.00653 | 0.00431 | [20] |
-Te iii | 5s2 5p3 | 0.08868 | 0.06478 | [10] | Pt i | 5d9 6s | 0.00543 | 0.00148 | [7] |
-Pt ii | 5d9 | 0.00444 | 0.00174 | [7] | |||||
--shell | Pt iii | 5d8 | 0.00309 | 0.00147 | [7] | ||||
-Fe ii | 3d6 4s | 0.00233 | 0.00058 | [12] | Au i | 5d10 6s | 0.01183 | 0.00554 | [21] |
-Fe iii | 3d6 | 0.00274 | 0.00075 | [13] | Au ii | 5d10 | 0.00471 | 0.00055 | [21] |
-Au iii | 5d9 | 0.00373 | 0.00176 | [21] |
+| -shell | 2.4 | -shell | continued | ||||||
+| Sr i | 4p6 5s2 | 5.4 | 2.6 | [9] | Ni ii | 3d9 | 2.9 | [14] | |
+| Y ii | 4p6 5s2 | 4.5 | 3.2 | [8] | Ni iii | 3d | 6.7 | [15] | |
+| Sr ii | 4p6 5s | 4.8 | 2.1 | [8] | Ni iv | 3d | 4.7 | [16] | |
+| Y iii | 4p6 4d | 3.7 | 2.3 | [11] | Zr i | 4d 5s2 | 5.0 | [17] | |
+| Zr ii | 4d2 5s | 1.7 | [17] | ||||||
+| -shell | 7.7 | Zr iii | 4d2 | 5.0 | [17] | ||||
+| Sr iii | 4p6 | 7.7 | 4.2 | [9] | W i | 6s2 5d4 | 2.8 | [18] | |
+| Te i | 5s2 5p4 | 6.7 | 1.2 | [10] | W ii | 6s2 5d3 | 3.7 | [19] | |
+| Te ii | 5s2 5p3 | 7.5 | 5.8 | [10] | W iii | 6s2 5d2 | 6.5 | [20] | |
+| Te iii | 5s2 5p2 | 8.9 | 6.5 | [10] | Pt i | 5d9 6s | 5.4 | [7] | |
+| Pt ii | 5d9 | 4.4 | [7] | ||||||
+| -shell | 8.1 | Pt iii | 5d8 | 3.1 | [7] | ||||
+| Fe ii | 3d6 4s | 2.3 | 5.8 | [12] | Au i | 5d10 6s | 1.2 | [21] | |
+| Fe iii | 3d6 | 2.7 | 7.5 | [13] | Au ii | 5d10 | 4.7 | [21] | |
+| Au iii | 5d9 | 3.7 | [21] |
 The emission spectrum as a function of wavelength in nm (scaled relative to the strongest line using only -matrix data) for Y ii is shown on Figure 2.
@@
 For this reason the optimized Axelrod coefficient is considerably higher than the surrounding -shell elements presented in Table 1.
-Figure 3 therefore demonstrates a mixed bag of agreement in the emission estimations because the Axelrod formula assumes that all the forbidden lines behave the same, which is evidently not the case for this ion.
+Figure 3 therefore demonstrates a lack of consistency in terms agreement in the emission estimations because the Axelrod formula assumes that all the forbidden lines behave the same, which is evidently not the case for this ion.
 It can be seen that certain lines are well represented with the R-matrix + Axelrod datasets (RM+AX), regardless of the Axelrod coefficient used.
@@
 Moving to the re-optimized Axelrod approximation introduces a certain improvement in the estimation of the emission, in many cases by an order of magnitude.
-The main victory is the improvement of the evidently strongest lines likely to contribute to a real astrophysical spectrum.
+The primary result is the improvement of the evidently strongest lines which are most likely to contribute to a real astrophysical spectrum.
 In principle we could optimise only on the strongest lines, although the optimization would depend on the plasma conditions chosen.
@@
 The elements studied cover the 1st (Sr-Zr), 2nd (Te) and 3rd (W,Pt,Au) peaks of the rapid nucleosynthesis processes.
-The values presented here could in principle be used to generate collision strengths for collisions between high-lying levels not included in structure calculations, or aid in the calculation of similarly structured elements by using ball-park coefficients reflective of those presented in Table 1.
+The values presented here could in principle be used to generate collision strengths for collisions between high-lying levels not included in structure calculations, or aid in the calculation of similarly structured elements by using approximate coefficients reflective of those presented in Table 1.
 Elements with -shell ground states tend to have transitions 10 - 1000 times stronger than the base Axelrod approximation.
@@
 Despite this, the relatively large variance in indicates that the Axelrod approximation should be reserved for where no data is available, and should be replaced with -matrix data as it comes available.
-Nonetheless, it is intended that this work serves as an intermediate step in the improvement of atomic data for astrophysical models, where the values presented here can be used to provide more information for elements with no data available and the Axelrod formula can be scaled based on the electronic configurations of the species involved.
+Nonetheless, it is intended that this work serves as an intermediate step in the improvement of atomic data for astrophysical models, where the values (or perhaps their mean values) presented here can be used to provide more information for elements with no data available and the Axelrod formula can be scaled based on the electronic configurations of the species involved.
 CRediT authorship contribution statement